A $25 000, 10% bond redeemable at par on December 1, 2025, is purchased on September 25, 2014, to yield 7.6% compounded semi-annually. Bond interest is payable semi-annually. Note: June 1, 2014 to December 1, 2025 = 11.5 years.

1.What is the cash price?

2. What is the accrued interest?

3.What is the quoted price?
Answer #1

1: Cash price=29546.64

Using financial calculator
Input: FV=25000

PMT=10%*25000/2 = 1250

I/Y=7.6/2 = 3.8

N=11.5*2 = 23

Solve for PV as -29546.64

2: Accrued interest = Coupon rate*FV*days till last interest/Total days

= 10%*25000*116/365

=794.52

3: Quoted price= 29546.64-794.52 = $ 28752.12